How much of you is not vacuum? Outside of the nucleus, an atom contains mostly empty space. Estimate the fraction of the volume of an atom that is occupied by the nucleus. Assume that a typical atom has a linear dimension of about 0.1 nm and that a typical nuclear dimension is about 1 fm.
1
Expert's answer
2020-04-30T10:42:00-0400
"\frac{V_n}{V_a}=\frac{a_n^3}{a_a^3}"

"\frac{V_n}{V_a}=\frac{(10^{-15})^3}{(0.1\cdot 10^{-9})^3}=10^{-15}"